OpenStudy (anonymous):

Harry is building a house with a sloping roof. If the roof is 44 feet long and 10 feet high, what is the angle that the roof makes with the horizontal?

OpenStudy (anonymous):

A. 12.8° B. 13.1° C. 14.6° D . 15.3° E. 17.5° which one is it

OpenStudy (anonymous):



OpenStudy (anonymous):

sin\(\theta\)=10/44 find \(\theta\)

OpenStudy (anonymous):

it will be 13.1

OpenStudy (anonymous):

oh thank so much